OpenStudy (anonymous):

2x+8 is less than or equal to x+2 is less than or equal to 3x+16 means that x is in the closed interval [a,b] where a: b:

OpenStudy (anonymous):

Are you able to solve\[2x+8 \le x+2\]

OpenStudy (anonymous):

2x+8=<x+2+=<3x+16

OpenStudy (hysusonic):

\[2x+8 \le x+2 \le 3x+16\]

OpenStudy (anonymous):

Consider it as two problems. Solving \[2x+8 \le x+2\] will give you b. Solving \[x+2 \le 3x+16\] will give you a.